How many grams of Cu would have to react to produce 89.5 grams of Ag?


1
Expert's answer
2022-02-19T06:57:47-0500

The balanced chemical equation:


Cu + 2AgNO3= Cu(NO3)2 + 2Ag


According to the equation: n(Cu) = n(AgNO3)/2

Moles of AgNO3 = Mass of AgNO3 / Molar mass of AgNO3

The molar mass of AgNO3 is 169.87 g mol

n(AgNO3) = Moles of AgNO3 = 45.0 g / 169.87 g mol-1 = 0.2649 mol

n(Cu) = n(AgNO3)/2 = 0.2649 mol / 2 = 0.13245 mol

Moles of Cu = Mass of Cu / Molar mass of Cu

The molar mass of Cu is 63.546 g mol-1.

Hence,

Mass of Cu = Moles of Cu × Molar mass of Cu = 0.13245 mol × 63.546 g mol-1 = 8.41667 g = 8.42 g

Mass of Cu is 8.42 grams.